OpenStudy (anonymous):

Reflect back on this scenario and each equation you created. Would any restrictions apply to the domain and range of those equations? Explain your reasoning using complete sentences.

OpenStudy (anonymous):

the equatons created are...

OpenStudy (anonymous):

x-y=-1 y-4=1(x-3) y=x+1 y-3=1(x-2) y=x+1